OpenStudy (anonymous):

Which of the following equations is parallel to the line passing through A(-2,7), B(-2,-2)?

OpenStudy (anonymous):

A – y=-2 B – y=x C – 3x = 5 D - -2x = 7y

OpenStudy (texaschic101):

we will find the slope first by using the slope formula : slope(m) = (y2 - y1) / (x2 - x1) (-2,7) x1 = -2 and y1 = 7 (-2,-2) x2 = -2 and y2 = -2 now lets sub slope(m) = (-2 - 7) / (-2 - (-2) m = -9/(-2 + 2) m = -9/0 The slope of this line is undefined....it is a vertical line. Since you are looking for a parallel line, the slope will stay the same. A vertical line is set up where x = ?. Your equation parallel to that line would be x = -2

OpenStudy (texaschic101):

However, that is not an answer choice......hmmm....I don't know then

OpenStudy (ranga):

Without much computation one can see the x values of both the points A and B are the same. That means the equation of that line is x = -2 and it is parallel to the y axis. Among the choices, which one has the same type of equation of x = constant? The third one has -3x = 5 Divide both sides by -3 x = -5/3 (which is a constant) So C is the answer.